OpenStudy (anonymous):

HHEEELLPPP!!! ;) what are the solutions of 4x^2-3x-5=0

OpenStudy (asnaseer):

what methods have you been taught?

OpenStudy (anonymous):

how can be equal 0 ?

OpenStudy (asnaseer):

you can use completing the square, or the formula for the solution of a quadratic equation

OpenStudy (anonymous):

.... SHOW ME!

OpenStudy (anonymous):

your equation is already written in standard form: \[\large ax^2+bx+c=0 \] where a, b, c are the coefficients in your equation. can you tell me what they are? a= b= c= ???

OpenStudy (anonymous):

HUH??

OpenStudy (anonymous):

match up your equation with this to identify a, b, c: \[\large ax^2+bx+c=0 \] \[\large 4x^2-3x-5=0 \] a= b= c= ???